OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6

OpenStudy (blurbendy):

use the formula: y - y1 = slope(x - x1) substitute x1 and y1 with your point (-2,-3)

OpenStudy (blurbendy):

understand?

OpenStudy (anonymous):

no @blurbendy

OpenStudy (blurbendy):

ok: y - (-3) = -6(x - (-2)) y + 3 = -6(x + 2) y + 3 = -6x - 12 y = -6x - 15

OpenStudy (anonymous):

so thats the answer ? @blurbendy

OpenStudy (blurbendy):

yes, do you understand how i got it? i just substitute (-2) in for x1 and (-3) in for y1 then solved to get it in the form y = mx + b

OpenStudy (anonymous):

but that isnt one of the choices @blurbendy

OpenStudy (zzr0ck3r):

we are given slope, so y = mx+b ----> y=-6x+b now use the point (-2,-3) to find the value of b -3=-6(-2)+b -3=12+b b = -15 so y=-6x-15

OpenStudy (zzr0ck3r):

this is correct, what are the options?

OpenStudy (anonymous):

y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (blurbendy):

it's the first one then. they just didnt put it in the form y = mx + b

OpenStudy (zzr0ck3r):

option 1 y + 3 = -6(x + 2) y+3 = -6x-12 y = -6x-15